# Durian: Dual Reference-guided Portrait Animation with Attribute Transferthanks: Project page: https://hyunsoocha.github.io/durian

### v1

@@ 4 Experiments
Dataset.
We train our model on the CelebV-Text (Yu et al., 2023), VFHQ (Xie et al., 2022), and Nersemble (Kirschstein et al., 2023) datasets, which together contain 2,747 videos. For evaluation, we randomly select 200 videos from VHFQ and CelebV-Text, including unseen identities, facial poses, and expressions. Since triplet datasets containing combinations of target attribute images, original portrait images, and edited portrait images are not available, we conduct a quantitative evaluation under a self-attribute transfer setting following HairFusion (Chung et al., 2025). Specifically, we sample two frames from each video as the portrait and attribute image, and generate animations conditioned on the facial keypoints. The outputs are then compared to groundtruth frames.
Metrics
Figure 4: Qualitative Comparison. We compare our method and the baselines that combine X-Portrait Xie et al. (2024) with face editing methods Chung et al. (2025); Yang et al. (2023); Zhang et al. (2025); Bilecen et al. (2024).
We evaluate each method using standard metrics for fidelity and perceptual quality. To assess reconstruction fidelity, we report $\text{L}_{1}$ distance, PSNR, SSIM, and LPIPS between generated frames and groundtruth. For perceptual realism, we compute the Fréchet Inception Distance (FID) (Parmar et al., 2022).
4.1 Comparison
Baselines.
As no prior work directly tackles portrait animation with attribute transfer from an in-the-wild reference, we construct two-stage baselines by combining image-level attribute transfer with video animation methods, resulting in 12 model combinations. For attribute transfer (stage 1), we consider: Paint-by-Example (PbE)Yang et al. (2023), a mask-conditioned diffusion method for reference image insertion; HairFusionChung et al. (2025) and StableHair (Zhang et al., 2025), diffusion-based models for hairstyle transfer with and without masks; and TriplaneEdit (Bilecen et al., 2024), a 3D-aware GAN-based face editor. For portrait animation (stage 2), we use: LivePortrait (Guo et al., 2024) (GAN-based keypoint animation), X-Portrait (Xie et al., 2024) (ReferenceNet + ControlNet), and MegActor-$\sum$ (Yang et al., 2025) (DiT-based video diffusion).
Results.
As shown in Table˜1, our method consistently outperforms all baseline combinations across both fidelity and perceptual quality metrics. Fig.˜4 presents a qualitative comparison against baselines using X-Portrait (Xie et al., 2024) as the animation module (stage 2). Our method generates coherent and realistic hairstyle animations that preserve the identity and maintain consistency in spatial extent, shape, and fine details across frames. Refer to the supplementary material for additional qualitative comparisons with other baseline combinations.
4.2 Ablation Study
Figure 5: Ablation Study. Removing each component and varying the training strategy results in visible degradation of synthesis quality.
@@ 4.2 Ablation Study
We evaluate the contributions of key components in our model and training strategy. Table˜2 presents quantitative results, and Fig.˜5 shows corresponding qualitative comparisons. “single ReferenceNet” replaces the dual-branch architecture with a shared encoder that receives the portrait and attribute images concatenated along the channel dimension, following CAT-VTON (Chong et al., 2024). This setup fails to separate the roles of the two inputs, resulting in undesired blending of attribute and identity cues. “w/o mask expansion” omits the attribute-aware augmentation that simulates variations in spatial extent. Without this strategy, the model tends to rely on the default shape of the portrait’s original attribute mask, making it less capable of handling diverse attribute shapes during inference. “w/o ref. image aug.” disables spatial and photometric augmentations applied to the reference images during training. As a result, the model fails to accurately transfer the desired attribute with misaligned reference images. “w/o ref. mask input” removes the binary mask concatenation from the inputs to the ReferenceNets. This weakens spatial localization and often leads to artifacts or residual traces of the original attribute in the output. “full ref. image input” uses unmasked portrait and attribute images during training. Interestingly, this variant achieves the best quantitative scores in Table˜2, which evaluates the self-attribute transfer setting, since full images simplify the task by allowing the model to copy content more easily. However, as shown in Fig.˜5, this model fails to disentangle identity and attribute roles, leading to visible identity leakage during cross-identity transfer. Ours achieves spatially consistent, identity-preserving results, and outperforms all other ablated variants except the full reference image variant.
@@ 4.3 Application
Figure 6: Multi-Attribute Transfer. Our model supports composition of multiple attributes (e.g., hair, eyeglasses, beard, hat) in a single forward pass.
Our model supports the composition of multiple attributes (e.g., glasses, hat, hairstyle) in a single generation pass by extending the spatial attention mechanism as described in Eq.˜17. Fig.˜6 show qualitative results where multiple attributes are simultaneously transferred from different reference images. Remarkably, our model not only combines multiple attributes seamlessly but also handles interactions between overlapping regions, such as between hair and a hat. Despite the reference images exhibiting diverse lighting conditions and spatial alignments, the model successfully integrates all attributes into the portrait image while maintaining a coherent and natural appearance.
@@
Our model supports the transfer of interpolated attributes between two different attribute images by linearly blending their reference features, as described in Eq.˜19. Fig.˜7 presents results for hair, beard, and eyeglasses, showing continuous transitions in both shape and appearance. The interpolations exhibit smooth changes in visual attributes such as color, volume, and style, demonstrating that our model effectively captures semantically meaningful directions in the attribute feature space.
@@ 5 Discussion
Figure 7: Attribute Interpolation. Our model enables smooth and consistent transitions between attributes. We present interpolation examples for hair, beard, and eyeglasses by varying the interpolation parameter $\alpha$.
@@
We present Durian, a zero-shot framework for portrait animation with cross-identity attribute transfer. Our method generates realistic, temporally coherent videos conditioned on a portrait image and a reference image depicting the target attribute. By leveraging self-reconstruction training on in-the-wild video frames, along with our attribute-aware mask expansion and augmentation strategy, our model learns spatially consistent and semantically grounded attribute transfer without requiring triplet supervision. Our model can also be extended to multi-attribute composition and attribute interpolation in a single generation pass without additional training.
However, several limitations remain. First, while our model supports multi-attribute transfer, complex interactions, such as hair occluded by hats, may not always be handled faithfully. Second, although we apply color-based augmentation during training, large lighting discrepancies between the portrait and attribute images can still degrade visual quality, particularly under strong shadows or highlights. Third, our training data primarily contains front-facing portraits; hence, transferring attributes from or to extreme poses (e.g., back views) remains an open challenge. Finally, because our method relies on accurate keypoint extraction and face tracking, failures in these components can lead to temporal jitter or artifacts during animation. Extending our approach to full-body scenarios and broader categories of attributes, such as clothing or accessories, presents an exciting direction for future work.
Appendix A Guidance Generation
@@
Our model generates portrait animations by leveraging a guidance video composed of facial keypoints, as shown in figure 2 of our main paper. These keypoints encode entangled facial shape information such as the distance between the eyes and ears, the interocular distance, and the distance from the eyes to the nose. While this rich representation contributes to accurate portrait animation in self-reconstruction scenarios, we observe that the animation tends to follow the facial shape of the guidance video in cross-identity reenactment scenarios. To address this, we propose a method that preserves the facial shape of the portrait image while still transferring the facial motion from a cross-identity guidance video. Specifically, we employ LivePortrait (Guo et al., 2024) to generate an animation video of the portrait image driven by the original guidance video from a different identity. We then extract the facial keypoint guidance video from this animation using Sapiens (Khirodkar et al., 2024) facial keypoint estimation.
Appendix B Additional Application
Text-to-Image-to-Attribute Transfer for Portrait Animation
Our method generates attribute transferred portrait animation video when an image containing the desired attribute is provided. We extend this capability by generating the attribute image from a text prompt, allowing for attribute-transferred portrait animation driven by textual descriptions, as shown in figure 8. We utilize the FLUX (Labs, 2024) text-to-image model to synthesize attribute images, which are then transferred to the portrait image to generate portrait animations.
Appendix C Additional Qualitative Comparison
@@
Extending the comparison presented in figure 4 of the main paper, we additionally provide qualitative results with other baseline combinations. Note that we generate portraits with transferred hair attributes using recent image insertion and face editing methods (Chung et al., 2025; Yang et al., 2023; Zhang et al., 2025; Bilecen et al., 2024), and compare the resulting animation videos produced by applying recent animation techniques (Guo et al., 2024; Xie et al., 2024; Yang et al., 2025) with those generated by our method, as shown in figure 9.
Figure 8: Text-to-Image-to-Attribute Transfer for Portrait Animation. We demonstrate portrait animation with attribute transfer from a textual description. We employ FLUX Labs (2024) to generate a high-quality portrait image with the desired hair attribute from a text prompt.
Figure 9: Qualitative Comparison. We compare our method and the baselines that combine portrait animation method  Guo et al. (2024); Xie et al. (2024); Yang et al. (2025) with face editing methods Chung et al. (2025); Yang et al. (2023); Zhang et al. (2025); Bilecen et al. (2024).
@@
Figure 10: More Qualitative Results. We present additional results on hair, hat, eyeglasses, and beard attribute transfer for portrait animation. Our method preserves the fine details of the original portrait while achieving natural and seamless attribute transfer.
@@
Figure 11: More Multi-Attribute Transfer Results. We demonstrate the results of simultaneously transferring two attributes for portrait animation.
@@
Figure 12: More Multi-Attribute Transfer Results. We present results of simultaneously transferring three attributes. In each example, the image at the top-left corner indicates the portrait that receives the attribute transfer.

### v2

@@ 4.2 Ablation Study
Figure 5: Ablation Study. Omitting components or altering training scheme degrades visual quality.
@@
We evaluate the contributions of key components in our model and training strategy. Table˜2 presents quantitative results, and Fig.˜5 shows corresponding qualitative comparisons. “single ReferenceNet” replaces the dual-branch architecture with a shared encoder that receives the portrait and attribute images concatenated along the channel dimension, following CAT-VTON (Chong et al., 2024). This setup fails to separate the roles of the two inputs, resulting in undesired blending of attribute and identity cues. “w/o mask expansion” omits the attribute-aware augmentation that simulates variations in spatial extent. Without this strategy, the model tends to rely on the default shape of the portrait’s original attribute mask, making it less capable of handling diverse attribute shapes during inference. “w/o ref. image aug.” disables spatial and photometric augmentations applied to the reference images during training. As a result, the model fails to accurately transfer the desired attribute with misaligned reference images. “w/o ref. mask input” removes the binary mask concatenation from the inputs to the ReferenceNets. This weakens spatial localization and often leads to artifacts or residual traces of the original attribute in the output. “full ref. image input” uses unmasked portrait and attribute images during training. Interestingly, this variant achieves the best quantitative scores in Table˜2, which evaluates the self-attribute transfer setting, since full images simplify the task by allowing the model to copy content more easily. However, as shown in Fig.˜5, this model fails to disentangle identity and attribute roles, leading to visible identity leakage during cross-identity transfer. Ours achieves spatially consistent, identity-preserving results, and quantitatively outperforms all other ablated variants except the full reference image variant.
@@ 4.3 Application
Figure 6: Multi-Attribute Transfer. Our model supports composition of multiple attributes (e.g., hair, eyeglasses, beard, hat) in a single forward pass without additional training.
Our model supports the composition of multiple attributes (e.g., glasses, hat, hairstyle) in a single generation pass by extending the spatial attention mechanism as described in Eq.˜12. Fig.˜6 show qualitative results where multiple attributes are simultaneously transferred from different reference images. Remarkably, our model not only combines multiple attributes seamlessly but also handles interactions between overlapping regions, such as between hair and a hat. Despite the reference images exhibiting diverse lighting conditions and spatial alignments, the model successfully integrates all attributes into the portrait image while maintaining a coherent and natural appearance.
@@
Our model enables attribute interpolation by linearly blending the reference features of two attributes, as in Eq.˜14. Fig.˜7 shows hair results with smooth transitions in shape and appearance. The interpolations exhibit smooth changes in visual attributes, demonstrating that our model effectively captures semantically meaningful directions in the attribute feature space.
@@ 5 Discussion
Figure 7: Attribute Interpolation. Our model enables smooth and consistent transitions between hair attributes by varying the interpolation parameter $\alpha$. More examples are in our Supp. Mat.
@@
We present Durian, a zero-shot framework for portrait animation with cross-identity attribute transfer, given a portrait image and one or more reference images specifying the target attributes. Our diffusion model, equipped with a Dual ReferenceNet, learns attribute transfer directly from uncurated portrait videos through a self-reconstruction training strategy, eliminating the need for triplet supervision. This is further enhanced by our attribute-aware mask expansion and augmentation scheme. Moreover, Durian naturally extends to multi-attribute composition and attribute interpolation within a single generation pass, without requiring any additional training.
Appendix A Implementation Details
A.1 Training Details
We adopt the two-stage training strategy following  Zhu et al. (2024). In the first stage, we resize all videos to a uniform resolution of $512\times 512$ pixels and train with a global batch size of 8 for 60,000 steps. During this phase, all layers except the temporal attention layers are set to be trainable, as the latter are not yet incorporated into the UNet. In the second stage, we insert temporal attention layers into the Denoising UNet (DNet) and train only these newly added layers. This stage uses 24-frame inputs, a global batch size of 8, and also runs for 60,000 steps. For both stages, we fix the learning rate at 1e-5, with each stage requiring approximately three days of training. We train our model using 8 NVIDIA RTX A6000 GPUs. As initialization, we use the UNet checkpoint from  Yang et al. (2023), while the temporal attention layers are initialized from  Guo et al. (2023). Our training dataset consists of 2,747 samples drawn from CelebV-Text (Yu et al., 2023), VFHQ (Xie et al., 2022), and Nersemble (Kirschstein et al., 2023). Our method focuses on four attribute categories, with the following distribution: Hair – 886 samples from CelebV-Text, 935 from Nersemble, and 265 from VFHQ (total 2,086); Beard – 253 samples from CelebV-Text; Eyeglasses – 279 samples from CelebV-Text; Hat – 129 samples from CelebV-Text. On average, each video contains 292 frames.
A.2 Evaluation Details
For self-attribute transfer, we randomly sample 200 videos from CelebV-Text (Yu et al., 2023) and VFHQ (Xie et al., 2022), ensuring that these videos contain unseen identities, facial poses, and expressions relative to the training dataset. For cross-attribute transfer, we additionally sample 50 videos. Masks required for image editing baselines are constructed following the procedures provided by the respective authors. To construct cross-attribute transfer pairs, we use the 50 sampled identities and randomly select corresponding face images from VFHQ and CelebV-Text that do not overlap with the training dataset.
We evaluate the results using several metrics. mCLIP-I (masked CLIP-I (Radford et al., 2021; Hessel et al., 2021)) and mDINO (Caron et al., 2021) (masked DINO) assess whether the target attribute is accurately transferred into the generated portrait animation video. To this end, we fill the background of attribute-only images with white and segment the target attribute region from the generated portrait animation video using Sapiens (Khirodkar et al., 2024). We then fill the segmented background with white and compute frame-wise cosine similarity embeddings with CLIP-I and DINO. ID-Sim evaluates identity preservation. Specifically, we mask attribute regions in portrait images by filling them with black, segment the target attribute regions in the generated videos with Sapiens, and replace them with black before computing frame-wise cosine similarity embeddings with ArcFace. Finally, VFID (Video Fréchet Inception Distance) (Heusel et al., 2017; Wang et al., 2018) extends FID to the video domain. Following  Fang et al. (2024), we adopt VFID to measure temporal consistency and overall video quality.
A.3 Keypoint guidance generation
@@
Our model generates portrait animations using a guidance video composed of facial keypoints, as shown in Fig.˜2 of our main paper. These keypoints encode entangled facial shape information, such as interocular distance and the relative positions of eyes, nose, and ears. While this rich representation supports accurate animation in self-attribute transfer scenarios, we observe that, in cross-attribute settings, the generated animation tends to follow the facial shape of the guidance video rather than the portrait image. To address this, we propose a method that preserves the portrait’s facial shape while transferring only the motion from a different identity. Specifically, we employ LivePortrait (Guo et al., 2024) to generate an animation of the portrait image that maintains its original shape while being driven by the motion in the guidance video. We then extract a facial keypoint guidance video from this animation using Sapiens (Khirodkar et al., 2024), effectively creating a self-reenactment-like scenario that allows our model to operate more reliably.
Appendix B Additional Results
B.1 Additional Ablation Study for Face Aligner
Figure 8: Ablation Study for Face Aligner. Omitting Face Aligner at inference time degrades the visual quality of the generated animation.
We perform an ablation study on our Face Aligner, as described in Section˜3.4 and illustrated in Fig.˜3 of the main paper. As shown in Fig.˜8, removing Face Aligner still allows the long blonde hair from the attribute image to be transferred to the portrait’s target attribute region. However, the generation becomes unstable, with the left hair strand intermittently appearing and disappearing. In contrast, ours, which applies the face aligner at inference time, enables stable transfer, ensuring that the long blond hair remains consistently preserved throughout the animation.
B.2 Additional Qualitative Comparison
Figure 9: Qualitative Comparison of Self-Attribute Transfer in the Hair Category. We compare our method and the baselines that combine portrait animation method with image or hairstyle editing methods. Our results show the highest quality closest to the ground truth, while other methods produce artifacts or unnatural appearances.
Qualitative comparison of self-attribute transfer.
@@ B.2 Additional Qualitative Comparison
We additionally provide qualitative results with other baseline combinations in a self-attribute transfer setup. Note that we generate portraits with transferred hair attributes using recent image insertion and face editing methods (Chung et al., 2025; Yang et al., 2023; Zhang et al., 2025; Bilecen et al., 2024), and compare the resulting animation videos produced by applying recent animation techniques (Guo et al., 2024; Xie et al., 2024; Yang et al., 2025) with those generated by our method, as shown in Fig.˜9.
Figure 10: Qualitative Comparison of Cross-Attribute Transfer in the Hair Category. We compare our method with the baselines that combine image editing and portrait animation. Our results best preserve the identity of the portrait image while most effectively transferring the hairstyle.
Qualitative comparison of cross-attribute transfer.
We extend the comparison in Fig.˜4 of the main paper and present results in Fig.˜10 against 12 baselines for cross-attribute transfer setup. Our method best preserves the identity of the portrait image while most accurately transferring the hairstyle from the attribute image. Furthermore, our results are perceived as the most natural and visually coherent.
B.3 Additional Quantitative Comparison
Figure 11: Qualitative Comparison of Self-Attribute Transfer in the Eyeglasses Category. TE represents TriplaneEdit and LP denotes LivePortrait. In the self-attribute transfer setting on the eyeglasses category, we compare our results with baseline. Our method produces portrait animations most similar to the ground truth while remaining the most natural.
Table 3: Quantitative Comparison on Eyeglasses Category. Our method outperforms this baseline on every evaluation metric.
[…]
TriplaneEdit LivePortrait 0.151 13.53 0.433 0.435 106.28
Ours 0.078 18.19 0.627 0.181 75.59
As shown in Fig.˜11 and Table˜3, we compare TriplaneEdit (Bilecen et al., 2024)+LivePortrait (Guo et al., 2024) with our method, since TriplaneEdit also supports transfer for eyeglasses. Our method consistently outperforms the baseline across all self-attribute transfer metrics. Moreover, it produces results that are closer to the ground truth and more natural than the baseline.
B.4 User Study
Table 4: User Study. We conduct a user study on two baseline methods that achieve strong performance in both self-attribute transfer and cross-attribute transfer. Our approach receives the highest preference among participants.
[…]
TriplaneEdit LivePortrait 4.45
PbE LivePortrait 19.04
Ours 76.50
We conduct a user study to evaluate portrait animations generated using portrait and attribute inputs from different identities, as shown in Table˜4. Each of the 100 participants viewed 9 randomly selected videos from a pool of 44 and rated how well each output preserved the hairstyle of the attribute image and the identity of the portrait image. Our method achieves the highest user preference, demonstrating superior performance in cross-identity transfer. Participants were asked: “Which video most naturally combines the face from the ‘face’ image with the hairstyle from the ‘hair’ image?”
B.5 Additional Results
@@
Figure 12: Qualitative Results for Single-Attribute Transfer. We present additional results on hair, hat, eyeglasses, and beard attribute transfer for portrait animation. Our method preserves the fine details of the original portrait while achieving natural and seamless attribute transfer.
Single-attribute transfer.
We extend the results of Fig.˜1 in the main paper and present in Fig.˜12 animations generated by transferring a single attribute to the portrait. Our method preserves the identity of the portrait image while faithfully transferring the attribute from the attribute image, resulting in natural portrait animations with attribute transfer.
@@
Figure 13: Qualitative Results for Dual-Attribute Transfer. We demonstrate the results of simultaneously transferring two attributes for portrait animation.
@@
Figure 14: Qualitative Results for Triple-Attribute Transfer. We present the results of simultaneously transferring three attributes. In each example, the image in the top-left corner indicates the target portrait.
Multi-attribute transfer.
In Fig.˜13 and Fig.˜14, we present portrait animations generated by simultaneously transferring two and three attributes in a single stage under the zero-shot setting. Through various combinations of the four supported categories (beard, eyeglasses, hair, hat), our method produces portrait animations where attributes are transferred naturally and with high quality, without any additional optimization.
Figure 15: Attribute Interpolation. We demonstrate smooth and consistent interpolation of additional attributes such as beard, eyeglasses, and hat according to the $\alpha$ values, extending beyond the hair interpolation results shown in the main paper.
Attribute interpolation.
We extend the results of Fig.˜7 in the main paper and present additional attribute interpolation results in Fig.˜15. Our method generates zero-shot, single-stage portrait animations with interpolated attributes, even for rigid objects such as hats and eyeglasses. The animations interpolate naturally according to the $\alpha$ values.
B.6 Text-to-Image Generated Attribute Transfer for Portrait Animation
Figure 16: Text-to-Image Generated Attribute Transfer for Portrait Animation. We generate a portrait animation with attribute transfer from a textual description by using FLUX (Labs, 2024) to synthesize a high-quality portrait image with the desired hair attribute.
Our method generates a portrait animation video with attribute transfer given an image containing the desired attribute. We extend this capability by synthesizing the attribute image directly from a text prompt, enabling text-driven control over the target attribute, as illustrated in Fig.˜16. Specifically, we leverage the FLUX (Labs, 2024) text-to-image model to generate realistic attribute images, which are then transferred to the portrait image to produce the final attribute-transferred portrait animation.